OpenStudy (anonymous):

If two adults and three children go to the movies, it costs $31.50. If one adult and five children go to the movies, it costs $35.00. What is the cost of one adult movie ticket?

OpenStudy (queelius):

2(adult cost) + 3(child cost) = 31.5 1(adult cost) + 5(child cost) = 35 Solve the system of linear equations.

OpenStudy (anonymous):

Say X is adult cost, Y = child cost. 2X+3Y=31.5 1X+5Y=35 Solve this system of equations to find the answer. after calculation, adult ticket: X= 5.5